# A Simple and Effective Reinforcement Learning Method for Text-to-Image Diffusion Fine-tuning

### v4

@@ 4 Method: Leave-One-Out PPO (LOOP) for Diffusion Fine-tuning
The expectation in the PPO loss (Eq. 6) is typically estimated by sampling a single trajectory for a given prompt $c$:
$\mbox{}\sum_{t=0}^{T}\text{clip}\left(\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})},1-\epsilon,1+\epsilon\!\right)r(\mathbf{x}_{0},\mathbf{c}),\!$ (13)
where $\mathbf{x}_{0:T}\sim\pi_{old}$. Even though the single sample estimate is an unbiased Monte-Carlo approximation of the expectation, it has high-variance (Owen, 2013). Additionally, the IS term ($\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}$) can also contribute to high-variance of the PPO objective (Swaminathan & Joachims, 2015; Xie et al., 2023). Both factors combined, can lead to high-variance, and unstable training of the PPO.
Taking inspiration from RLOO (Eq. 10), we sample $K$ independent trajectories from the previous policy for a given prompt $c$, and apply a baseline correction term from each trajectories reward, to reduce the variance of the estimator:
$\begin{split}&\hat{J}_{\theta}^{\mathrm{LOOP}}(\pi)={}\\ &\frac{1}{K}\!\sum_{i=1}^{K}\!\bigg{[}\!\sum_{t=0}^{T}\text{clip}\!\left(\frac{\pi_{\theta}(\mathbf{x}^{i}_{t-1}|\mathbf{x}^{i}_{t},c)}{\pi_{\text{old}}(\mathbf{x}^{i}_{t-1}|\mathbf{x}^{i}_{t},c)},1-\epsilon,1+\epsilon\right)\\ &\mbox{}\hskip 130.88284pt\cdot\left(r(\mathbf{x}^{i}_{0},\mathbf{c})-b^{i}\right)\Big{]},\end{split}$ (14)
where $\mathbf{x}^{i}_{0:T}\sim\pi_{old},\forall i\in[1,K]$. The baseline correction term $b^{i}$ reduces the variance of the gradient estimate, while being unbiased in expectation (Gupta et al., 2024a; Mohamed et al., 2020). A simple choice of baseline correction can be the average reward across the $K$ trajectories, i.e.:
$b^{i}=\frac{1}{k}\sum_{i=0}^{K}r(\mathbf{x}^{i}_{0}).$ (15)
However, we choose the leave-one-out average baseline, with average taken across all samples in the trajectory, except the current sample $i$, i.e.:
$b^{i}=\frac{1}{k-1}\sum_{j\neq i}r(\mathbf{x}^{j}_{0}).$ (16)
Originally RLOO sampling and baseline corrections were proposed in the context of REINFORCE, with a focus on on-policy optimization (Ahmadian et al., 2024; Kool et al., 2019), whereas we are applying these in the off-policy step of PPO. We call this method leave-one-out PPO (LOOP). Provenly, LOOP has lower variance than PPO:
Proposition 4.1.
The LOOP estimator $\hat{J}_{\theta}^{\mathrm{LOOP}}(\pi)$ (Eq. 14) has lower variance than the PPO estimator $\hat{J}_{\theta}^{\mathrm{PPO}}(\pi)$ (Eq. 13):
$\mathrm{Var}\left[\hat{J}_{\theta}^{\mathrm{LOOP}}(\pi)\right]<\mathrm{Var}\left[\hat{J}_{\theta}^{\mathrm{PPO}}(\pi)\right].$ (17)
Proof.
Since the sampled trajectories are independent:
$\mathrm{Var}\!\left[\hat{J}_{\theta}^{\mathrm{LOOP}}\!(\pi)\right]\!=\!\frac{1}{K^{2}}\mathrm{Var}\!\left[\hat{J}_{\theta}^{\mathrm{PPO}}\!(\pi)\right]\!<\!\mathrm{Var}\!\left[\hat{J}_{\theta}^{\mathrm{PPO}}\!(\pi)\right].\quad\;\qed$ (18)
5 Experimental Setup
Stable v1.4 (Rombach et al., 2022) 0.3765 0.3576 0.4156 0.1246 0.4461
Stable v2 (Rombach et al., 2022) 0.5065 0.4221 0.4922 0.1342 0.4579
Composable v2 (Liu et al., 2022) 0.4063 0.3299 0.3645 0.0800 0.4261
Structured v2 (Feng et al., 2022) 0.4990 0.4218 0.4900 0.1386 0.4550
Attn-Exct v2 (Chefer et al., 2023) 0.6400 0.4517 0.5963 0.1455 0.4767
GORS unbiased (Huang et al., 2023) 0.6414 0.4546 0.6025 0.1725
GORS (Huang et al., 2023) 0.6603 0.4785 0.6287 0.1815 0.4841
PPO (Black et al., 2023) 0.6821 0.5655 0.6909 0.1961 0.5102
LOOP ($k=2$) 0.6785 0.5746 0.6937 0.1800 0.5072
LOOP ($k=3$) 0.7515 0.6220 0.7353 0.1966 0.5242
LOOP ($k=4$) 0.7859 0.6676 0.7518 0.2136 0.5422
Benchmark. Text-to-image diffusion and language models often fail to satisfy an essential reasoning skill of attribute binding. Attribute binding reasoning capability refers to the ability of a model to generate images with attributes such as color, shape, texture, spatial alignment, (and others) specified in the input prompt. In other words, generated images often fail to bind certain attributes specified in the instruction prompt (Huang et al., 2023; Ramesh et al., 2022; Fu & Cheng, 2024).
Since attribute binding seems to be a basic requirement for useful real-world applications, we choose the T2I-CompBench benchmark (Huang et al., 2023), which contains multiple attribute binding/image compositionality tasks, and its corresponding reward metric to benchmark text-to-image generative models. We also select two common tasks from RL for diffusion works: improving aesthetic quality of generation, and image-text semantic alignment (Black et al., 2023; Fan et al., 2024). To summarize, we choose the following tasks for the RL optimization: (i) Color, (ii) Shape, (iii) Texture, (iv) 2D Spatial, (v) Numeracy, (vi) Aesthetic, (vii) Image-text Alignment. For all tasks, the prompts are split into training/validation prompts. We report the average reward on both training and validation split.
Model. As the base diffusion model, we use Stable diffusion V2 (Rombach et al., 2022), which is a latent diffusion model. For optimization, we fully update the UNet model, with a learning rate of $1e^{-5}$. We also tried LORA fine-tuning (Hu et al., 2021), but the results were not satisfactory, so we update the entire model instead. The hyper-parameters are reported in Appendix A.
6 Results and Discussion
@@
Table 2: Comparing REINFORCE with PPO on the T2I-CompBench benchmark over three image attributes: Color, Shape, and Texture. The metrics in this table are average reward on an unseen test set (higher is better). For each prompt, average rewards over 10 independent generated images are calculated.
@@
PPO 0.6821 0.5655 0.6909
6.1 REINFORCE vs. PPO efficiency-effectiveness trade- off
We discuss our empirical results for the REINFORCE vs. PPO efficiency-effectiveness trade-off. Our empirical validation of the trade-off compares the following methods:
REINFORCE. The REINFORCE policy gradient for diffusion fine-tuning (Eq. 8).
REINFORCE with baseline correction. We compare the REINFORCE policy gradient with a baseline correction (BC) term (Eq. 9). For the baseline term, we choose the average reward for the given prompt (Black et al., 2023).
PPO. The PPO objective for diffusion fine-tuning with importance sampling and clipping (Eq. 6).
Figure 2 shows the training reward over epochs for the attributes: Color, Shape, and Texture from the T2I-CompBench benchmark, and training reward from optimizing the aesthetic model. It is clear that REINFORCE policy gradient is not effective in terms of performance, as compared to other variants. Adding a baseline correction term indeed improves the training performance, validating the effectiveness of baseline in terms of training performance, possibly because of reduced variance. PPO achieves the highest training reward, validating the effectiveness of importance sampling and clipping for diffusion fine-tuning.
Table 3: Comparing the performance of LOOP with PPO on the aesthetic and image-text alignment tasks. Higher values are better.
@@
Color Shape Texture Numeracy
Reward
Refer to caption Refer to caption Refer to caption Refer to caption
      Spatial       Image-Text Alignment       Aesthetic
Reward
Refer to caption Refer to caption Refer to caption
Figure 3: Comparing PPO with proposed LOOP on the T2I-CompBench benchmark w.r.t. image attributes: Color, Shape, Texture, Numeracy, and Spatial relationship. We also compare against aesthetic and image-text alignment tasks (Black et al., 2023). The y-axis corresponds to the training reward; the x-axis corresponds to the training epoch. Results are averaged over 3 independent training runs; shaded areas indicate 80% prediction intervals.
      Color       Aesthetic
Variance
Refer to caption Refer to caption
       Epochs        Epochs
Figure 4: Training reward variance for the color, and aesthetic task. The y-axis corresponds to the training reward variance; the x-axis corresponds to the training epoch. Results are averaged over 3 runs; shaded areas indicate 80% prediction intervals. We observe that higher values of samples reused (i.e., $k$) produce lower reward variance during training.
We also evaluate the performance on a separate validation set. For each validation prompt, we generate 10 independent images from the diffusion policy, and average the reward, finally averaging over all evaluation prompts. The validation results are reported in Table 2. The results are consistent with the pattern observed with the training rewards, i.e., REINFORCE with baseline provides a better performance than plain REINFORCE, suggesting that baseline correction indeed helps with the final performance. Nevertheless, PPO still performs better than REINFORCE.
6.2 Evaluating LOOP
Performance during training. Figure 3 shows the training reward curves for different tasks, against number of epochs. LOOP outperforms PPO across all seven tasks consistently throughout training. This establishes the effectiveness of sampling multiple diffusion trajectories per input prompt, and the leave-one-out baseline correction term (Eq. 10) during training. Training reward curve is smoother for the aesthetic task, as compared to tasks from the T2I-CompBench benchmark. We hypothesise that improving the attribute binding property of diffusion model is a harder task than improving the aesthetic quality of generated images.
Table 1 reports the validation rewards across different tasks from the T2I-CompBench benchmark. LOOP outperforms PPO and other strong supervised learning based baseline significantly across all tasks. It shows that PPO improves the attribute-binding reasoning ability of the diffusion model compared to other supervised learning based methods.
For the aesthetic and image-text alignment objectives, the validation rewards are reported in Table 3. LOOP results in a 15.37% relative improvement over PPO for the aesthetic task, and a 2.16% improvement over PPO for the image-text alignment task.
Impact of number of independent trajectories ($k$). The LOOP variant with number of independent trajectories $K=4$ performs the best across all tasks, followed by the variant $K=3$. This is intuitive given that Monte-Carlo estimates get better with more number of samples (Owen, 2013). Surprisingly, the performance of the variant with $K=2$ is comparable to PPO.
Impact on training variance. We evaluate whether LOOP results in a lower empirical variance than PPO, as proved theoretically in Lemma 4.1. Figure 4 reports the empirical reward variance during training for the color attribute and aesthetic objective. LOOP results in a lower empirical variance than PPO, thereby empirically validating our claim that LOOP has lower variance than PPO.
Qualitative results. For a qualitative evaluation of the attribute-binding reasoning ability, we present some example image generations from SD, PPO, and LOOP in Figure 1. In the first example, the input prompt specifies a black colored ball with a white cat. Stable diffusion (SD) and PPO fail to bind the color black with the generated ball, whereas LOOP successfully binds that attribute. Similarly, in the third example, SD and PPO fail to bind the hexagon shape attribute to the watermelon, whereas LOOP manages to do that. In the fourth example, SD and PPO fail to add the horse object itself, whereas LOOP adds the horse with the specified black color, and flowing cyan patterns.
7 Conclusion
We have studied the efficiency-effectiveness trade-off between two fundamental RL for diffusion methods: REINFORCE, and PPO. REINFORCE, while computationally efficient and easier to implement, is subpar to PPO in terms of sample efficiency and performance. Building on these insights, we have introduced a simple and effective RL for diffusion method, LOOP, which builds on the variance reduction techniques from REINFORCE and the effectiveness and robustness of PPO. We have found that LOOP improves over diffusion models on multiple black-box objectives. A limitation of LOOP is that sampling multiple diffusion trajectories per prompt can lead to more computational overhead and an increase in training time. A potential future direction would be to keep the effectiveness of LOOP while maintaining the computational complexity of PPO.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
Appendix A Hyperparameter and Implementation Details
For REINFORCE (including REINFORCE with baseline correction term), PPO, and LOOP the number of denoising steps ($T$) is set to 50. The diffusion guidance weight is set to 5.0. For optimization, we use AdamW (Loshchilov & Hutter, 2017) with a learning rate of $1e^{-5}$, and the weight decay of $1e{-4}$, with other parameters kept at the default value. We clip the gradient norm to $1.0$. We train all models using 8 A100 GPUs with a batch size of 4 per GPU. The clipping parameter $\epsilon$ for PPO, and LOOP is set to $1e^{-4}$.
Appendix B Additional Qualitative Examples
We present some additional qualitative examples in this section.

### v5

@@ 4 Method: Leave-One-Out PPO (LOOP) for Diffusion Fine-tuning
The expectation in the PPO loss (Eq. 3) is typically estimated by sampling a single trajectory from the policy in the previous iteration $\pi_{old}$: $\mathbf{x}_{0:T}\sim\pi_{old}$ for a given prompt $c$:
$\mbox{}\hskip-2.84526pt\sum_{t=0}^{T}\text{clip}\left(\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})},1-\epsilon,1+\epsilon\!\right)r(\mathbf{x}_{0},\mathbf{c}).\!$ (12)
Even though the single sample estimate is an unbiased Monte-Carlo approximation of the expectation, it has high-variance (mcbook). Additionally, the IS term ($\frac{\pi_{\theta}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}{\pi_{\text{old}}(\mathbf{x}_{t-1}|\mathbf{x}_{t},\mathbf{c})}$) can also contribute to high-variance of the PPO objective (swaminathan2015self; xie2023dropout). Both factors combined, can lead to high-variance, and unstable training of the PPO.
Taking inspiration from RLOO (Eq. 9), we sample $K$ independent trajectories from the previous policy for a given prompt $c$, and apply a baseline correction term from each trajectory’s reward, to reduce the variance of the estimator:
$\hat{J}_{\theta}^{\mathrm{LOOP}}(\pi)=\frac{1}{K}\!\sum_{i=1}^{K}\!\bigg[\!\sum_{t=0}^{T}\text{clip}\!\left(\frac{\pi_{\theta}(\mathbf{x}^{i}_{t-1}|\mathbf{x}^{i}_{t},c)}{\pi_{\text{old}}(\mathbf{x}^{i}_{t-1}|\mathbf{x}^{i}_{t},c)},1-\epsilon,1+\epsilon\right)\cdot\left(r(\mathbf{x}^{i}_{0},\mathbf{c})-b^{i}\right)\Big],$ (13)
where $\mathbf{x}^{i}_{0:T}\sim\pi_{old},\forall i\in[1,K]$. The baseline correction term $b^{i}$ reduces the variance of the gradient estimate, while being unbiased in expectation (gupta2024optimal; mohamed2020monte). A simple choice of baseline correction can be the average reward across the $K$ trajectories. However, it results in a biased estimator (kool2019buy). Therefore, we choose the leave-one-out average baseline, with average taken across all samples in the trajectory, except the current sample $i$, i.e.:
$b^{i}=\frac{1}{k-1}\sum_{j\neq i}r(\mathbf{x}^{j}_{0}).$ (14)
Originally RLOO sampling and baseline corrections were proposed in the context of REINFORCE, with a focus on on-policy optimization (ahmadian2024back; kool2019buy), whereas we are applying these in the off-policy step of PPO. We call this method leave-one-out PPO (LOOP).
Our approach is conceptually similar to the recently popular GRPO method for RL fine-tuning of LLMs (shao2024deepseekmath). Although our work was developed independently before GRPO gained widespread recognition, we do not include a head-to-head comparison.
Technically, the distinction lies in following aspects: (i) unlike GRPO, our formulation does not apply standard-deviation normalization in the denominator, as this has been shown to potentially harm performance in recent LLM fine-tuning via RL studies (liu2025understanding), (ii) similar to GRPO, we omit a KL penalty term, since our empirical experiments showed that it has little practical benefit. Furthermore, a recent study showed that on-policy RL implicitly constrains the updated policy to remain close to the base policy under a KL divergence measure, even without an explicit KL penalty term (shenfeld2025rl), and (iii) we ignore the generation-length normalization term. In the diffusion setting, this simplification is further justified by the fact that the sequence length of the reverse diffusion process is fixed across generations, rendering length normalization unnecessary.
5 Experimental Setup
Benchmark. Text-to-image diffusion and language models often fail to satisfy an essential reasoning skill of attribute binding. Attribute binding reasoning capability refers to the ability of a model to generate images with attributes such as color, shape, texture, spatial alignment, (and others) specified in the input prompt. In other words, generated images often fail to bind certain attributes specified in the instruction prompt (huang2023t2i; ramesh2022hierarchical; fu2024enhancing). Since attribute binding seems to be a basic requirement for useful real-world applications, we choose the T2I-CompBench benchmark (huang2023t2i), which contains multiple attribute binding/image compositionality tasks, and its corresponding reward metric to benchmark text-to-image generative models. We also select two common tasks from prior RL for diffusion work: improving aesthetic quality of generation, and image-text semantic alignment (black2023training; fan2024reinforcement). To summarize, we choose the following tasks for the RL optimization: (i) Color, (ii) Shape, (iii) Texture, (iv) 2D Spatial, (v) Numeracy, (vi) Aesthetic, (vii) Image-text Alignment. For all tasks, the prompts are split into training/validation prompts. We report the average reward on both training and validation split.
Model. As the base diffusion model, we use Stable diffusion V2 (rombach2022high), which is a latent diffusion model. For optimization, we fully update the UNet model, with a learning rate of $1e^{-5}$. We also tried LORA fine-tuning (hu2021lora), but the results were not satisfactory, so we update the entire model instead.
6 Hyperparameter and Implementation Details
For REINFORCE (including REINFORCE with baseline correction term), PPO, and LOOP the number of denoising steps ($T$) is set to 50. The diffusion guidance weight is set to 5.0. For optimization, we use AdamW Loshchilov2017DecoupledWD with a learning rate of $1e^{-5}$, and the weight decay of 1e41e{-4}, with other parameters kept at the default value. We clip the gradient norm to $1.0$. We train all models using 8 A100 GPUs with a batch size of 4 per GPU. The clipping parameter $\epsilon$ for PPO, and LOOP is set to $1e^{-4}$.
7 Results and Discussion
7.1 REINFORCE vs. PPO efficiency-effectiveness trade- off
We present our empirical results on the efficiency-effectiveness trade-off between REINFORCE and PPO. Our evaluation compares the following methods: the REINFORCE policy gradient for diffusion fine-tuning (Eq. 7); the REINFORCE policy gradient with a baseline correction term (REINFORCE w/ BC), detailed in Eq. 8, where the baseline term is the average reward for the given prompt (black2023training), and the PPO objective for diffusion fine-tuning, which incorporates importance sampling and clipping, as outlined in Eq. 3. This PPO objective is equivalent to the DDPO objective in the original RL for diffusion method (black2023training).
Figure 2 shows the training reward over epochs for the attributes: Color, Shape, and Texture from the T2I-CompBench benchmark, and training reward from optimizing the aesthetic model. Results are averaged over 3 runs. It is clear that REINFORCE policy gradient is not effective in terms of performance, as compared to other variants. Adding a baseline correction term indeed improves the training performance, validating the effectiveness of baseline in terms of training performance, possibly because of reduced variance. PPO achieves the highest training reward, validating the effectiveness of importance sampling and clipping for diffusion fine-tuning.
@@ 7 Results and Discussion
We also evaluate the performance on a separate validation set. For each validation prompt, we generate 10 independent images from the diffusion policy, and average the reward, finally averaging over all evaluation prompts. The validation results are reported in Table 1. The results are consistent with the pattern observed with the training rewards, i.e., REINFORCE with baseline provides a better performance than plain REINFORCE, suggesting that baseline correction indeed helps with the final performance. Nevertheless, PPO (DDPO) still performs better than REINFORCE.
@@
Table 1: Comparing REINFORCE with DDPO on the T2I-CompBench benchmark over three image attributes: Color, Shape, and Texture. We report average reward on unseen test set (higher is better). For each prompt, average rewards over 10 independent generated images are calculated.
@@
DDPO 0.6821 0.5655 0.6909
7.2 Evaluating LOOP
Performance during training. Figure 3 shows the training reward curves for different tasks, against number of epochs. LOOP outperforms DDPO (black2023training) across all seven tasks consistently throughout training. This establishes the effectiveness of sampling multiple diffusion trajectories per input prompt, and the leave-one-out baseline correction term (Eq. 9) during training. Training reward curve is smoother for the aesthetic task, as compared to tasks from the T2I-CompBench benchmark. We hypothesise that improving the attribute binding property of diffusion model is a harder task than improving the aesthetic quality of generated images.
Color Shape Texture
Reward
Refer to caption Refer to caption Refer to caption
    Spatial     Image-Text Alignment     Aesthetic
Reward
Refer to caption Refer to caption Refer to caption
Figure 3: Comparing DDPO (referenced as PPO) with the proposed LOOP on the T2I-CompBench benchmark with respect to image attributes: Color, Shape, Texture, and Spatial relationship. We also report results on aesthetic preference and image–text alignment tasks (black2023training). The y-axis shows training reward, and the x-axis shows training epoch. Results are averaged over three independent runs; shaded areas denote 80% prediction intervals.
Stable v1.4 (rombach2022high) 0.3765 0.3576 0.4156 0.1246 0.4461
Stable v2 (rombach2022high) 0.5065 0.4221 0.4922 0.1342 0.4579
Composable v2 (liu2022compositional) 0.4063 0.3299 0.3645 0.0800 0.4261
Structured v2 (feng2022training) 0.4990 0.4218 0.4900 0.1386 0.4550
Attn-Exct v2 (chefer2023attend) 0.6400 0.4517 0.5963 0.1455 0.4767
GORS unbiased (huang2023t2i) 0.6414 0.4546 0.6025 0.1725
GORS (huang2023t2i) 0.6603 0.4785 0.6287 0.1815 0.4841
DDPO (black2023training) 0.6821 0.5655 0.6909 0.1961 0.5102
LOOP ($k=3$) 0.7515 0.6220 0.7353 0.1966 0.5242
LOOP ($k=4$) 0.7859 0.6676 0.7518 0.2136 0.5422
Table 2 reports the average rewards on the test set across various tasks from the T2I-CompBench benchmark. For each prompt, we generate 10 different images and calculate the average rewards. LOOP consistently outperforms DDPO (black2023training) and other strong supervised learning-based baselines across all tasks. Notably, LOOP achieves relative improvements of 18.1% and 15.2% over DDPO on shape and color attributes, respectively.
For the aesthetic and image-text alignment objectives, the validation rewards are reported in Table 3. LOOP results in a 15.4% relative improvement over PPO for the aesthetic task, and a 2.4% improvement over PPO for the image-text alignment task.
Impact of number of independent trajectories ($k$). The LOOP variant with number of independent trajectories $K=4$ performs the best across all tasks, followed by the variant $K=3$. This is intuitive given that Monte-Carlo estimates get better with more number of samples (mcbook). Surprisingly, the performance of the variant with $K=2$ is comparable to PPO.
8 Qualitative Examples
For a qualitative evaluation of the attribute-binding reasoning ability, we present some example image generations from SD, DDPO, and LOOP in Figures 1, 4, and 5.
In Figure 1 qualitative examples of the attribute binding task are presented. In the example in the first column of Figure 1, input prompt specifies a black ball with a white cat. Stable diffusion (SD) and PPO fail to bind the color black with the generated ball, whereas LOOP successfully binds that attribute. Similarly, in the third column, SD and PPO fail to bind the hexagon shape attribute to the watermelon, whereas LOOP manages to do that. In the fourth column, SD and PPO fail to add the horse object itself, whereas LOOP adds the horse with the specified black color, and flowing cyan patterns.
Figure 4 highlights improvements in aesthetic quality of the generated images. Compared to SD v2 and PPO, LOOP produces sharper, more coherent compositions with balanced lighting and color tone. For example, in the second column (“a cat”) and in the fourth column (“butterfly”), LOOP enhances realism and contrast while preserving overall artistic intent.
Finally, Figure 5 presents additional qualitative examples that emphasize both binding and aesthetics. LOOP accurately binds challenging color-object pairs (e.g., teal branch, pink cornfield) while producing more visually appealing and natural results. PPO and SD v2 often miss attribute alignment or produce dull, less cohesive scenes.
Table 3: Comparing the performance of LOOP with DDPO on the aesthetic and image-text alignment tasks. Higher values are better.
@@ 8 Qualitative Examples
9 Conclusion
We have studied the efficiency-effectiveness trade-off between two fundamental RL methods for diffusion fine-tuning: REINFORCE and PPO. Our analysis, both theoretical and empirical, demonstrates that while REINFORCE is computationally efficient and easier to implement, it suffers from high variance and sample inefficiency compared to PPO. PPO, though more effective, comes with significant computational overhead, requiring three models in memory simultaneously and involving sensitive hyperparameter tuning.
Building on these insights, we have introduced LOOP, a novel RL method for diffusion fine-tuning that combines variance reduction techniques from REINFORCE (multiple trajectory sampling and leave-one-out baseline correction) with the robustness and sample efficiency of PPO (importance sampling and clipping). Our empirical evaluation on the T2I-CompBench benchmark demonstrates that LOOP achieves substantial improvements over both the base Stable Diffusion model and the state-of-the-art PPO method across multiple tasks, including attribute binding (color, shape, texture, spatial relationships), aesthetic quality, and image-text alignment.
Quantitatively, LOOP (k=4) achieves substantial improvements over PPO across all evaluated tasks. On the T2I-CompBench benchmark, LOOP achieves relative improvements of $\mathbf{18.1}\%$ on shape binding, $\mathbf{15.2}\%$ on color binding, $\mathbf{8.8}\%$ on texture binding, and $\mathbf{8.9}\%$ on spatial reasoning. LOOP also improves aesthetic quality by $\mathbf{15.4}\%$ and image-text alignment by $\mathbf{2.2}\%$. Qualitatively, as shown in Figures 1, 4, and 5, LOOP successfully binds attributes that previous methods fail to capture, while also producing more visually coherent and aesthetic images.
A limitation of LOOP is the increased computational cost from sampling multiple diffusion trajectories per prompt, which leads to longer training times compared to standard PPO. Future work could explore adaptive sampling strategies to reduce this overhead while maintaining LOOP’s effectiveness, extend the method to other diffusion architectures and modalities, or investigate the integration of human preference modeling for better alignment with real-world objectives.